# Multi-Scale Finetuning for Encoder-based Time Series Foundation Models

### v1

@@ Abstract
Time series foundation models (TSFMs) demonstrate impressive zero-shot performance for time series forecasting. However, an important yet underexplored challenge is how to effectively finetune TSFMs on specific downstream tasks. While naive finetuning can yield performance gains, we argue that it falls short of fully leveraging TSFMs’ capabilities, often resulting in overfitting and suboptimal performance. Given the diverse temporal patterns across sampling scales and the inherent multi-scale forecasting capabilities of TSFMs, we adopt a causal perspective to analyze finetuning process, through which we highlight the critical importance of explicitly modeling multiple scales and reveal the shortcomings of naive approaches. Focusing on encoder-based TSFMs, we propose Multiscale finetuning (MSFT), a simple yet general framework that explicitly integrates multi-scale modeling into the finetuning process. Experimental results on three different backbones (Moirai, Moment and UniTS) demonstrate that TSFMs finetuned with MSFT not only outperform naive and typical parameter efficient finetuning methods but also surpass state-of-the-art deep learning methods.
@@ 2 Preliminaries
In multi-scale modeling, the standard approach for generating multi-scale sequences is based on average pooling [35, 42]. Given a training sample $(\mathbf{X},\mathbf{Y})$, both context and horizon windows are downsampled into multiple temporal scales using non-overlapping average pooling. Specifically, downsampling factor is commonly set to 2, resulting in a set of scales defined by ${1,2,\dots,2^{K}}$, where $K$ is the number of downsampled scales. Let $\mathcal{S}$ denote the set of multi-scale time series as $\mathcal{S}=\{\mathbf{S}_{0},\dots,\mathbf{S}_{K}\}$, where $\mathbf{S}_{i}=(\mathbf{X}^{i},\mathbf{Y}^{i})$ corresponds to the $i$-th scale series, formed by concatenating the downsampled context $\mathbf{X}^{i}\in\mathbb{R}^{C_{i}\times D}$ and downsampled horizon $\mathbf{Y}^{i}\in\mathbb{R}^{H_{i}\times D}$. Here, $C_{i}=\lceil\frac{C}{2^{i}}\rceil$ and $H_{i}=\lceil\frac{H}{2^{i}}\rceil$. Note that $\mathbf{S}_{0}$ represents the input series at the original scale.
@@
[…]
@@ 3.1 Multi-Scale Effect on TSFM: A Causal View
It is evident that scale $S$ is a confounder that induces spurious correlations between input context series (via $S\rightarrow X$) and activated knowledge of TSFM (via $S\rightarrow M$). The former captures the multi-scale properties of time series, while the latter corresponds to the multi-scale capabilities of TSFM. Scale $S$ ultimately affects the forecasting of the prediction horizon via the backdoor path $X\leftarrow S\rightarrow M\rightarrow Y$. Naive finetuned forecaster for $P(Y|X)$ overlooks the impact of this backdoor path, learning forecasting only at the original scale. This oversight would mistakenly associate non-causal but positively correlated input context to forecast horizon in the original scale, resulting in problematic forecasting.
@@ 4 Methodology
To address the aforementioned challenges, we propose MSFT to realize the high-level framework in Figure 2(a) as an effective multiscale finetuning strategy. Specifically, to activate scale-specific knowledge, we freeze the pretrained parameters and introduce scale-specific, parameter-efficient modules into the 1  input projection and 2 attention layers. To eliminate the cross-scale attention bias and correctly capture temporal correlations, we propose a decoupled token dependency modeling mechanism: I  in-scale self-attention captures within-scale dependencies, while II cross-scale aggregators explicitly fuse information across scales, ensuring correct temporal alignment between tokens. Finally, we apply multi-scale mixing to the 3 output projection, combining scale-specific predictions with learned weights. Figure 3 illustrates our MSFT method.
@@
$\operatorname{C2F:}\quad$ ${\bm{h}}^{l}_{i-1}={\bm{h}}^{l}_{i-1}+\operatorname{Repeat}(\tilde{{\bm{h}}}^{l}_{i,i-1}),\quad$ for $\text{for }i\in\{K,...,1\}$ (3)
$\operatorname{F2C:}\quad$ ${\bm{h}}^{l}_{i+1}={\bm{h}}^{l}_{i+1}+\operatorname{AvgPool}(\tilde{{\bm{h}}}^{l}_{i,i+1}),\quad$ for $\text{for }i\in\{0,...,K-1\}$ (4)
@@ 6.3 Model Analysis
To fully understand MSFT, we conduct model analysis using the MoiraiSmall model on three LSF datasets, selected for its strong zero-shot performance and relatively low training cost. Due to page limits, we present the analysis of down-sampling approaches, down-sampling factors, detailed attention analysis, and visualizations in the Appendix C. We also discuss the potential application of MSFT to decoder-based structures and its limitation in Appendix D.
@@
Ablations 1 to 4 examine the effectiveness of scale-specific knowledge activation. For both input projection and attention, either freezing ( 1, 3) or finetuning shared weights ( 2, 4) yields inferior performance to using scale-specific modules, with freezing causing larger performance drops. Among the two, attention has a greater impact than input projection, highlighting its critical role in capturing temporal dependencies.
@@
Figure 5 shows the attention score heatmaps of three attention strategies. In (a), direct attention (Ablation 8) exhibits spurious temporal dependencies, with attention scores biased toward tokens sharing the same time indices. In (b), we align time indices during attention, ensuring that cross-scale tokens corresponding to the same temporal region share identical time indices. While this approach produces "correct" attention patterns, it is limited to RoPE and performs worse than our method (see Appendix C for details). In (c), our in-scale masking strategy eliminates misleading cross-scale attention, focusing on accurate within-scale dependency modeling.
@@ 7 Conclusion
This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
@@ A.2 Multi-scale modeling
Besides the down-sampling-based methods discussed in Section 5.2, there is another strategy multi-patch, which applies specifically to patch-based TSTs. Here, multiple patch sizes are used to segment the time series into tokens in different resolutions. Pathformer [6] applies layer-by-layer routing to select patch sizes and aggregate the outputs from multiple scales. MTST [49] proposes a multi-branch architecture for modeling diverse temporal patterns at different resolutions. ElasTST [47] leverages a shared transformer backbone with tunable RoPE for multi-scale patch assembly. However, this line of methods requires the model to be compatible with multiple patch sizes, which is not feasible for most TSFMs.
Despite these efforts to incorporate multi-scale features, our proposed method exhibit two key differences from current approaches. First, most of these models are trained from scratch on a single dataset, leaving open questions about how to effectively leverage pretrained TSFMs for multi-scale modeling. Our method is the first to explore multi-scale modeling directly on pretrained TSFMs. Second, existing methods primarily perform cross-scale aggregation at the feature or time series level, whereas our approach achieves this at the token level.
@@ Appendix B Implementation Details
B.1 Dataset details
@@
B.2 Encoder-based TSFMs
In this section, we describe the architectural details and training objectives of each encoder-based TSFM used in our experiments. Table 6 summarizes the fundamental characteristics of the models based on their origin setup.
@@ B.4 Metric details
[…]
@@ B.5 Experiment Details
Appendix C More Experimental Results
C.1 Further Model Analysis
@@
C.2 Full results

### v2

@@ Abstract
Time series foundation models (TSFMs) demonstrate impressive zero-shot performance for time series forecasting. However, an important yet underexplored challenge is how to effectively finetune TSFMs on specific downstream tasks. While naive finetuning can yield performance gains, we argue that it falls short of fully leveraging TSFMs’ capabilities, often resulting in overfitting and suboptimal performance. Given the diverse temporal patterns across sampling scales and the inherent multi-scale forecasting capabilities of TSFMs, we adopt a causal perspective to analyze finetuning process, through which we highlight the critical importance of explicitly modeling multiple scales and reveal the shortcomings of naive approaches. Focusing on encoder-based TSFMs, we propose Multiscale finetuning (MSFT), a simple yet general framework that explicitly integrates multi-scale modeling into the finetuning process. Experimental results on three different backbones (Moirai, Moment and UniTS) demonstrate that TSFMs finetuned with MSFT not only outperform naive and typical parameter efficient finetuning methods but also surpass state-of-the-art deep learning methods. Codes are available at https://github.com/zqiao11/MSFT.
@@ 2 Preliminaries
In multi-scale modeling (see Appendix A.2 for the detailed definition of this concept), the standard approach for generating multi-scale sequences is based on average pooling [35, 43]. Given a training sample $(\mathbf{X},\mathbf{Y})$, both context and horizon windows are downsampled into multiple temporal scales using non-overlapping average pooling. Specifically, downsampling factor is commonly set to 2, resulting in a set of scales defined by $1,2,\dots,2^{K}$}, where $K$ is the number of downsampled scales. Let $\mathcal{S}$ denote the set of multi-scale time series as $\mathcal{S}=\{\mathbf{S}_{0},\dots,\mathbf{S}_{K}\}$, where $\mathbf{S}_{i}=(\mathbf{X}^{i},\mathbf{Y}^{i})$ corresponds to the $i$-th scale series, formed by concatenating the downsampled context $\mathbf{X}^{i}\in\mathbb{R}^{C_{i}\times D}$ and downsampled horizon $\mathbf{Y}^{i}\in\mathbb{R}^{H_{i}\times D}$. Here, $C_{i}=\lceil\frac{C}{2^{i}}\rceil$ and $H_{i}=\lceil\frac{H}{2^{i}}\rceil$. Note that $\mathbf{S}_{0}$ represents the input series at the original scale.
@@
$\begin{aligned} {\bm{h}}^{0}&=\operatorname{InProject}({\bm{x}});&{\bm{h}}^{l}&=\operatorname{AttnBlock}({\bm{h}}^{l-1}),\,l=1,...,L;&\hat{\mathbf{Y}}&=\operatorname{OutProject}({\bm{h}}^{L})\end{aligned}$ (1)
@@ 3.1 Multi-Scale Effect on TSFM: A Causal View
It is evident that scale $S$ is a confounder that induces spurious correlations between input context series (via $S\rightarrow X$) and activated knowledge of TSFM (via $S\rightarrow M$). The former captures the multi-scale properties of time series, while the latter corresponds to the multi-scale capabilities of TSFM. Scale $S$ ultimately affects the forecasting of the prediction horizon via the backdoor path $X\leftarrow S\rightarrow M\rightarrow Y$. Naive finetuned forecaster for $P(Y|X)$ overlooks the impact of this backdoor path, learning forecasting only at the original scale. This oversight would mistakenly associate non-causal but positively correlated input context to forecast horizon in the original scale, resulting in problematic forecasting. Further discussion can be found in Appendix C.
@@ 4 Methodology
To address the aforementioned challenges, we propose MSFT to realize the high-level framework in Figure 2(a) as an effective multiscale finetuning strategy. Specifically, to activate scale-specific knowledge, we freeze the pretrained parameters and introduce scale-specific, parameter-efficient modules into the 1 input projection and 2 attention layers. To eliminate the cross-scale attention bias and correctly capture temporal correlations, we propose a decoupled token dependency modeling mechanism: I in-scale self-attention captures within-scale dependencies, while $I$ cross-scale aggregators explicitly fuse information across scales, ensuring correct temporal alignment between tokens. Finally, we apply multi-scale mixing to the 3 output projection, combining scale-specific predictions with learned weights. Figure 3 illustrates our MSFT method, with the detailed pseudo-code provided in Appendix B.1.
@@
$\operatorname{C2F:}\quad$ ${\bm{h}}^{l}_{i-1}={\bm{h}}^{l}_{i-1}+\operatorname{Repeat}(\tilde{{\bm{h}}}^{l}_{i,i-1}),\quad$ $\text{for }i\in\{K,...,1\}$ (3)
$\operatorname{F2C:}\quad$ ${\bm{h}}^{l}_{i+1}={\bm{h}}^{l}_{i+1}+\operatorname{AvgPool}(\tilde{{\bm{h}}}^{l}_{i,i+1}),\quad$ $\text{for }i\in\{0,...,K-1\}$ (4)
@@ 6.3 Model Analysis
To fully understand MSFT, we conduct model analysis using the MoiraiSmall model on three LSF datasets, selected for its strong zero-shot performance and relatively low training cost. Due to page limits, we present the analysis of down-sampling approaches, down-sampling factors, detailed attention analysis, and visualizations in the Appendix D. We also discuss the potential application of MSFT to decoder-based structures and its limitation in Appendix E.
@@
To ensure statistical robustness, we report mean results over three runs in Table 3, with standard deviations provided in Table 11. Ablations 1 to 4 examine the effectiveness of scale-specific knowledge activation. For both input projection and attention, either freezing ( 1, 3) or finetuning shared weights ( 2, 4) yields inferior performance to using scale-specific modules, with freezing causing larger performance drops. Among the two, attention has a greater impact than input projection, highlighting its critical role in capturing temporal dependencies.
@@
Figure 5 shows the attention score heatmaps of three attention strategies. In (a), direct attention (Ablation 8) exhibits spurious temporal dependencies, with attention scores biased toward tokens sharing the same time indices. In (b), we align time indices during attention, ensuring that cross-scale tokens corresponding to the same temporal region share identical time indices. While this approach produces "correct" attention patterns, it is limited to RoPE and performs worse than our method (see Appendix D for details). In (c), our in-scale masking strategy eliminates misleading cross-scale attention, focusing on accurate within-scale dependency modeling.
@@ 7 Conclusion
This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. Zhongzheng Qiao is affiliated with Energy Research Institute @ NTU, Interdisciplinary Graduate Programme, Nanyang Technological University, Singapore.
@@ A.2 Multi-scale modeling
The term multi-scale modeling has been used inconsistently in prior time series works. In our paper, it specifically refers to constructing multiple downsampled versions of the same time series and jointly leveraging them through cross-scale aggregation during prediction. This enables the model to integrate temporal information from both coarse and fine resolutions within a single forward pass.
Several related terminologies exist in the literature. TTM[11] introduces multi-resolution pretraining by downsampling high-frequency datasets into lower-frequency versions, which serves as a form of data augmentation during pretraining. However, each time series sample is still processed in a single scale/resolution during its prediction process. Thus, TTM does not perform multi-scale modeling as defined in our work. Another similar strategy is multi-patch-size modeling, which applies specifically to patch-based TSTs. Here, multiple patch sizes are used to segment the time series into tokens in different resolutions. Pathformer [6] applies layer-by-layer routing to select patch sizes and aggregate the outputs from multiple scales. MTST [50] proposes a multi-branch architecture for modeling diverse temporal patterns at different resolutions. ElasTST [48] leverages a shared transformer backbone with tunable RoPE for multi-scale patch assembly. Moirai [44] adopts multiple patch-size projection layers, yet it does not downsample inputs into multiple temporal scales. In summary, this line of methods requires the model to be compatible with multiple patch sizes, which is not feasible for most TSFMs. Under this clarified definition, our work is the first to introduce multi-scale modeling in TSFM finetuning.
@@ Appendix B Implementation Details
B.1 Pseudo-code of MSFT
For clarity, we provide the Pytorch-like pseudo codes of MSFT in Algorithm 1 and Algorithm 2, , illustrating the overall training pipeline and the MSFT attention block described in Section 4. Importantly, the method does not construct a new model from scratch, but enhances the pretrained TSFM through additional plug-in modules.
Algorithm 1 Overall Training Pipeline for MSFT: PyTorch-like Pseudo-code
class MSFTPipeline($n$.Module):
def __init__(self, pretrained_model, K, s=2, P=16):
self.K = K # number of new scales
self.s = s # downsample factor
self.P = P # patch size
# frozen input projection, scale-specific linear layers
self.in_proj = pretrained_model.in_proj
self.linears = [Linear_i() for i in range(K+1)]
# encoder layers with MSFT blocks
self.encoder = $n$.ModuleList([
AttnBlock_with_MSFT(block, K=K, s=s)
for block in pretrained_model.encoder_layers
# frozen output projection
self.out_proj = pretrained_model.out_proj
def forward(self, X, Y):
# Step 1: Multi-Scale Generation
S = []
for i in range(self.K+1):
X_i = AvgPool(X, window_size=self.s**i) # pre-pad if needed
Y_i = AvgPool(Y, window_size=self.s**i) # post-pad if needed
S.append((X_i, Y_i))
# Step 2: Patching & Projection
H_0 = []
for i, (X_i, Y_i) in enumerate(S):
x_i = Patching((X_i, Y_i), patch_size=self.P)
h_i = self.linears[i](self.in_proj(x_i)) # frozen InProject
h_i = Masking(h_i) # mask prediction tokens
H_0.append(h_i)
h = Concat(H_0)
scale_index = GetScaleIndex(H_0)
# Step 3: Multi-Scale Attention Encoding
for l in range(self.L):
h = self.encoder[l](h, scale_index)
# Step 4: Output & Loss
H_L = Split(h, scale_index) # recover [h_0^L, ..., h_K^L]
losses, preds = [], []
for i, (_, Y_i) in enumerate(S):
Y_hat_i = self.out_proj(H_L[i])
L_i = Loss(Y_i, Y_hat_i)
losses.append(w_i * L_i) # weighted by learnable w_i
# upsample prediction back to original scale
Y_hat_up = Upsample(Y_hat_i, scale=self.s**i)
preds.append(w_i * Y_hat_up)
L_total = sum(losses)
Y_hat = sum(preds)
return L_total, Y_hat
Algorithm 2 Multi-Scale Attention Block (AttnBlock_with_MSFT): PyTorch-like Pseudo-code
class AttnBlock_with_MSFT(PretrainedAttnBlock):
def __init__(self, K, s=2):
super().__init__()
self.K = K # number of new scales
self.s = s # downsample factor
# LoRA adapters for each scale
self.W_Q = [LoRA(self.W_Q) for _ in range(K+1)]
self.W_K = [LoRA(self.W_K) for _ in range(K+1)]
self.W_V = [LoRA(self.W_V) for _ in range(K+1)]
# cross-scale projections
self.F2CMap = [Linear() for i in range(K)] # fine -> coarse
self.C2FMap = [Linear() for i in range(K)] # coarse -> fine
def forward(self, h_in, scale_index):
# Step 1: Split input into scale-wise representation
H_in = [h_in[..., idx, :] for idx in scale_index]
# Step 2: Scale-specific Attention with LoRA
Q, K, V = [], [], []
for i in range(self.K+1):
Q.append(W_Q[i](H_in[i]))
K.append(W_K[i](H_in[i]))
V.append(W_V[i](H_in[i]))
Q, K, V = Concat(Q), Concat(K), Concat(V)
# Step 3: In-scale masked attention
h_attn = ScaledDotProductAttention(Q, K, V, mask=M_in)
# Step 4: Cross-scale Aggregation
H_attn = [h_attn[..., idx, :] for idx in scale_index]
# (a) Coarse-to-Fine (C2F)
H_c2f = H_attn.copy()
for i in range(self.K, 0, -1):
h_proj = self.C2FMap[i-1](H_attn[i])
H_c2f[i-1] += Repeat(h_proj, repeat_factor=self.s)
# (b) Fine-to-Coarse (F2C)
H_f2c = H_attn.copy()
for i in range(self.K):
h_proj = self.F2CMap[i](H_attn[i])
H_f2c[i+1] += AvgPool(h_proj, pool_size=self.s)
# (c) Merge outputs from two branches
H_out = []
for i in range(self.K+1):
H_out.append(0.5 * (H_c2f[i] + H_f2c[i]))
# Step 5: Re-concatenate
h_out = Concat(H_out)
# W_o & Add & Norm & FeedForward omitted for brevity
return h_out
B.2 Dataset details
@@
B.3 Encoder-based TSFMs
We describe the architectural details and training objectives of each encoder-based TSFM used in our experiments. Table 6 summarizes the fundamental details of the models based on their origin setup.
@@ B.5 Metric details
CRPS $=\int_{0}^{1}2\Lambda_{\alpha}(F^{-1}(\alpha),\mathbf{Y})d\alpha$
@@ B.6 Experiment Details
Appendix C Causal Analysis
C.1 Causal Modeling Motivation
Here we elaborate the motivation of our causal modeling in Figure 2. A time series can be viewed as a discretized sequence of sampled observations derived from an underlying continuous process. Under this perspective, the observed input window $X$ corresponds to a discrete observation of the latent process during a context period. The variable $X$ arises from two factors: the latent continuous process $I$ (unobserved) and the scale parameter $S$, which governs the sampling resolution. The scale $S$ determines how densely the latent process is sampled, thereby shaping both the temporal granularity and the length of the observed sequence $X$. While our formulation omits the latent process $I$ for tractability, the edge $S\to X$ in our causal graph reflects this observation mechanism. Importantly, scale influences the form of the observed input series but not the latent process itself.
Formally, we treat both $S$ and $X$ as random variables. The scale $S$ is a discrete variable that selects the resolution level for downsampling the input context. It takes values from a finite index set $\mathcal{S}=\{s_{0},s_{1},\ldots,s_{K}\}$, where each $s_{k}$ corresponds to a specific downsampling factor or temporal resolution. The observed input $X$ is then a random variable whose sequence length depends on the selected scale $S$. Thus, $X\in\mathcal{X}$, where $\mathcal{X}=\bigcup_{s\in\mathcal{S}}\mathbb{R}^{L_{s}}$ and $L_{s}$ denotes the input length corresponding to scale $s$. This formalization clarifies the role of scale in shaping observed input time series, consistent with the causal edge $S\to X$ in our proposed graph.
C.2 Empirical Validation of the Causal Graph
To empirically validate the proposed causal graph, we perform causal structure learning and partial correlation analysis on the ETTm1 dataset across multiple scales. Specifically, we extract context windows from the training split, downsample each window into three additional resolutions, and feed each scaled input into the pretrained Moirai model to obtain corresponding embeddings $\mathbf{M}$. For each sample at each scale, we form a triplet $(S,X,M)$, where $S$ is the scale index, $X$ is the autocorrelation (ACF) computed on the input, and $M$ is the $\ell_{2}$ norm of the mean embedding. These triplets enable graph-based causal discovery.
We first apply the PC algorithm [37, 51] with Fisher’s $Z$-test ($\alpha=0.01$). The learned graph includes directed edges $S\to X$ and $S\to M$, supporting our assumption that scale causally influences both the input signal and the model representation. To further test whether $S$ acts as a confounder between $X$ and $M$, we compare their raw correlation with the partial correlation conditioned on $S$. The Pearson correlation between $X$ and $M$ is 0.732-0.732; conditioning on $S$ reduces the partial correlation to 0.481-0.481 ($p<0.001$). This reduction indicates that scale partially explains the dependency between $X$ and $M$, consistent with its role as a confounder in our causal formulation. Together, these complementary analyses quantitatively support the causal assumption proposed in Section 3, namely that the scale variable $S$ influences both the observed input $X$ and the model knowledge $M$.
Appendix D More Experimental Results
D.1 Further Model Analysis
@@
We also compare the computational and parameter efficiency of MSFT with several representative fine-tuning strategies, including full finetuning, linear probing, LoRA, and AdaLoRA. Results on MoiraiSmallwith the Weather LSF task are summarized in Table 9.
Table 9: Quantitative comparison of computation efficiency across finetuning methods.
Method Params (M) GPU Mem (MB) Train Speed (it/s) Test Speed (it/s) MSE / MAE
Full finetuning 13.8 2996 9.4 104.5 0.228 / 0.254
Linear probing 3.0 762 19.7 113.6 0.237 / 0.260
LoRA 3.4 2760 11.1 115.9 0.225 / 0.252
AdaLoRA 3.4 2756 10.0 102.4 0.226 / 0.252
MSFT (ours) 4.4 5616 4.6 102.5 0.216 / 0.248
As shown in Table 9, MSFT uses fewer trainable parameters than full finetuning and achieves the best forecasting accuracy (lowest MSE/MAE). The cost is higher GPU memory usage and slower training due to the expanded token length introduced by multi-scale inputs. Nevertheless, its test-time speed remains comparable, offering a reasonable efficiency–performance trade-off compared to other fine-tuning methods.
More Ablation Study
To assess robustness, we conducted multiple runs (three seeds) for the experiments in Table 3. Results in Table 11 show that the standard deviations are generally insignificant across settings.
To further validate the role of in-scale masking, we conduct additional ablations where cross-scale aggregators are retained even without masking. Results in Table 10 show that this setting yields worse performance, as attention without in-scale masking learns misaligned cross-scale dependencies, and subsequent aggregation amplifies these inconsistencies. This confirms that cross-scale aggregators must operate jointly with in-scale masking to effectively fuse temporal correlations, while their standalone use significantly compromises performance.
Table 10: Ablation studies on in-scale masking, with mean $\pm$ standard deviation over 3 runs.
Config ETTm1 ETTm2 Weather
MSE MAE MSE MAE MSE MAE
8 0.360 $\pm$ 0.003 0.380 $\pm$ 0.002 0.253 $\pm$ 0.000 0.303 $\pm$ 0.000 0.220 $\pm$ 0.000 0.253 $\pm$ 0.001
8+ C2F 0.364 $\pm$ 0.004 0.383 $\pm$ 0.003 0.256 $\pm$ 0.002 0.307 $\pm$ 0.002 0.224 $\pm$ 0.000 0.256 $\pm$ 0.000
8+ F2C 0.365 $\pm$ 0.003 0.384 $\pm$ 0.002 0.253 $\pm$ 0.000 0.303 $\pm$ 0.001 0.225 $\pm$ 0.001 0.255 $\pm$ 0.001
8+ Both 0.365 $\pm$ 0.003 0.385 $\pm$ 0.003 0.255 $\pm$ 0.002 0.305 $\pm$ 0.001 0.226 $\pm$ 0.001 0.256 $\pm$ 0.001
MSFT 0.354 $\pm$ 0.003 0.378 $\pm$ 0.002 0.250 $\pm$ 0.000 0.301 $\pm$ 0.000 0.216 $\pm$ 0.000 0.248 $\pm$ 0.000
Table 11: Ablation study on three LSF datasets using MoiraiSmall. Mean $\pm$ standard deviation over 3 runs are reported. MSFT consistently outperforms all ablation variants.
Config ETTm1 ETTm2 Weather
MSE MAE MSE MAE MSE MAE
1 0.362 $\pm$ 0.003 0.380 $\pm$ 0.002 0.253 $\pm$ 0.001 0.305 $\pm$ 0.000 0.219 $\pm$ 0.000 0.252 $\pm$ 0.000
2 0.360 $\pm$ 0.002 0.379 $\pm$ 0.002 0.252 $\pm$ 0.000 0.304 $\pm$ 0.000 0.218 $\pm$ 0.000 0.249 $\pm$ 0.000
3 0.374 $\pm$ 0.004 0.385 $\pm$ 0.003 0.256 $\pm$ 0.001 0.308 $\pm$ 0.001 0.224 $\pm$ 0.002 0.256 $\pm$ 0.001
4 0.361 $\pm$ 0.002 0.382 $\pm$ 0.002 0.254 $\pm$ 0.000 0.306 $\pm$ 0.000 0.222 $\pm$ 0.001 0.254 $\pm$ 0.001
5 0.371 $\pm$ 0.003 0.384 $\pm$ 0.002 0.256 $\pm$ 0.000 0.307 $\pm$ 0.001 0.223 $\pm$ 0.001 0.255 $\pm$ 0.001
6 0.363 $\pm$ 0.003 0.382 $\pm$ 0.002 0.254 $\pm$ 0.000 0.304 $\pm$ 0.000 0.220 $\pm$ 0.000 0.252 $\pm$ 0.000
7 0.357 $\pm$ 0.002 0.379 $\pm$ 0.001 0.252 $\pm$ 0.000 0.304 $\pm$ 0.000 0.218 $\pm$ 0.000 0.251 $\pm$ 0.000
8 0.360 $\pm$ 0.003 0.380 $\pm$ 0.002 0.253 $\pm$ 0.000 0.303 $\pm$ 0.000 0.220 $\pm$ 0.001 0.253 $\pm$ 0.001
9 0.359 $\pm$ 0.003 0.379 $\pm$ 0.003 0.269 $\pm$ 0.003 0.313 $\pm$ 0.002 0.226 $\pm$ 0.003 0.252 $\pm$ 0.002
10 0.384 $\pm$ 0.007 0.388 $\pm$ 0.004 0.255 $\pm$ 0.003 0.311 $\pm$ 0.002 0.219 $\pm$ 0.001 0.252 $\pm$ 0.000
MSFT 0.354 $\pm$ 0.003 0.378 $\pm$ 0.002 0.250 $\pm$ 0.000 0.301 $\pm$ 0.000 0.216 $\pm$ 0.000 0.248 $\pm$ 0.000
D.2 Evaluation of knowledge forgetting
To further investigate the potential issue of knowledge forgetting, we conduct a simple zero-shot transfer experiment following the setup of [42]. Specifically, we finetune the model on a source dataset $A$ and directly evaluate it on an unseen target dataset $B$, denoted as $A\to B$. Table 12 reports MSE averaged over four prediction lengths for MoiraiSmall. The results reveal no consistent pattern across transfer settings. In some cases (e.g., ETTm1 $\to$ ETTm2), fine-tuning leads to improved zero-shot generalization, while in others (e.g., ETTm2 $\to$ ETTm1) the performance degrades, suggesting that finetuning overrides certain pretrained knowledge. When comparing full finetuning and MSFT, neither method consistently outperforms the other, indicating that both approaches are not explicitly designed to mitigate catastrophic forgetting. Moreover, the multi-scale knowledge learned by MSFT from dataset $A$ may not always generalize to dataset $B$ if their temporal structures differ significantly.
Table 12: Cross-dataset transfer performance (MSE) on MoiraiSmall. Each entry reports the performance when the model is finetuned on dataset $A$ and evaluated on an unseen dataset $B$ (denoted $A\to B$). The results are averaged over four prediction lengths. Best values are highlighted in bold.
Transfer Zero-shot Full FT MSFT
ETTm1 $\to$ ETTm2 0.300 0.293 0.288
ETTm2 $\to$ ETTm1 0.448 0.454 0.470
ETTm1 $\to$ ETTh1 0.416 0.410 0.420
ETTm2 $\to$ ETTh1 0.416 0.415 0.414
ETTm1 $\to$ ETTh2 0.355 0.350 0.363
ETTm2 $\to$ ETTh2 0.355 0.359 0.350
From another perspective, MSFT is inherently more conservative in overwriting pretrained representations, due to its plug-in design. During finetuning, only lightweight adapters, normalization layers, and the output head are updated, while the majority of pretrained weights remain frozen. This design principle is consistent with common strategies in continual learning, where task-specific modules are introduced to reduce forgetting [2]. Furthermore, if users wish to preserve the zero-shot performance on unseen datasets after finetuning, one can simply deactivate or remove the MSFT modules, effectively reverting the model to its original pretrained TSFM with minimal performance change.
D.3 Full results